OpenStudy (anonymous):

Find the inter-quartile range of the given data set. 3, 16, 18, 25, 26, 29, 31, 50

OpenStudy (anonymous):

q3 -q1

OpenStudy (anonymous):

ok first find the the ! it is the median of 3,16,18,25 Q1=16+18/2=34/2=17 now Q3 is the median of 26, 29, 31, 50 Q3=29+31/2=60/2=30 IQR=q3-Q1 can you find it now? after finding show me answer

OpenStudy (anonymous):

@samyflashy

OpenStudy (anonymous):

13

OpenStudy (anonymous):

yes you got it.its correct

OpenStudy (anonymous):

THANKS :)

OpenStudy (anonymous):

yw:)